If p = – 2, find the value of:
4p + 7 …..............

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that p = -2
Then, substitute the value of p in the question
= (4 × (-2)) + 7
= -8 + 7
= -1

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Given : p= -2
Substituting value of m in the equation,
4p + 7
= (4×-2) + 7
= -8+7
= -1
 
Hope it helps!
Thank you ✌️